a plane can fly 700 miles in the same time a car to drive 250 miles. how fast are the car and plane moving if the plane flies 90 miles per hour faster than the car
Answer by lwsshak3(11628)

a plane can fly 700 miles in the same time a car to drive 250 miles. how fast are the car and plane moving if the plane flies 90 miles per hour faster than the car
..
let x = speed of car
then, x+90 = speed of plane
travel time = distance/speed
travel time of car = 250/x
travel time of plane = 700/x+90
travel time of car=travel time of plane
250/x=700/90+x
700x=250x+22500
450x=22500
x=50 mph (speed of car)
x+90=140 mph (speed of plane)